When 7 is added to the reciprocal of twice a number, the result is the reciprocal of 4 times the number. Find the number?
Answer by mananth(16946)

let number be x
(1/2x) +7 =1/4x
(1/2x)-(1/4x)=-7
LCM = 4x
(2-1)/4x=-7
1=28x
x=-1/28